OpenStudy (anonymous):

A progression has a first term of 12 and a fifth term of 18. Find the sum of the first 25 terms in the progression is aritmetic.

OpenStudy (anonymous):

Is the difference 1.5?

OpenStudy (callisto):

+d +d +d +d 12 -->a2 ---> a3 --->a4 --->18 therefore, 12+4d = 18, d= (18-12)/4 = 1.5

OpenStudy (anonymous):

Ok, so what's the sum?

OpenStudy (callisto):

T(25) = a +(n-1)(1.5) = 12+(25-1)(1.5) = 48 Sum = (12+48)*25/2 = 1500/2 =750

OpenStudy (anonymous):

Thank you :)

OpenStudy (callisto):

you're welcome. I don't know if i get it correct :S

OpenStudy (anonymous):

I tried and got the same answer. Can you see the most recent post I posted?